OpenStudy (anonymous):

Solve: negative 1 over 2 x + 2 = −x + 7

OpenStudy (lollylau):

-1/2x+2=-x+7 -1+4x=-2x^2+14x 2x^2-10x-1=0 "confused"

OpenStudy (anonymous):

kk